# Inferring Foresightedness in Dynamic Noncooperative Games

### v2

@@ V-B Optimizing Parameters with Gradient Descent
The only non-trivial term in Equation 11 is $\nabla_{(\theta,\gamma)}v$. Next, we show how to take directional derivatives of Equation 4 at a solution $v^{*}$ where strictly complementarity holds.
First, we must consider the complementarity constraints on $z$ and $h(\cdot)$. To this end, we construct an index set $\mathcal{I}$ which records all inactive inequality constraint dimensions in $h(v;\theta,\gamma)$. Indexing $F$ at elements of this set yields a vector $[F]_{\mathcal{I}}$. From Equation 4b, we know that these inactive inequalities $[F]_{\mathcal{I}}$ are strictly positive, and the Lagrange multipliers associated with these constraints are exactly 00. Then, presuming the continuity of $F$, small changes in $(\theta,\gamma)$ preserve the positivity of elements of $[F]_{\mathcal{I}}$ and force the corresponding Lagrange multipliers in $[v]_{\mathcal{I}}$ to remain 00. Thus, we find that $\nabla_{(\theta,\gamma)}[v]_{\mathcal{I}}=0$.
Consider the remaining constraints from Equation 8, which must be active due to strict complementarity; denote the indices corresponding to these constraint as $\mathcal{S}$. We can use the implicit function theorem and stationarity of $F$ with respect to $v$ to write
00 0 $=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}+(\nabla_{v}[F]_{\mathcal{S}})(\nabla_{(\theta,\gamma)}v)$ (12a)
@@
Then, when $\nabla_{v}[F]_{\mathcal{S}}$ is invertible, we can find exact values of $\nabla_{(\theta,\gamma)}v$. For a more complete treatment, including a discussion of weak complementarity, readers are encouraged to consult [dontchev2009implicit, facchinei_pang, liu2023learningplaytrajectorygames].
Thus equipped, we iteratively update our estimate of $(\theta,\gamma)$ according to the gradient of Equation 6a—incorporating the aforementioned implicit derivatives as needed. Algorithm 1 terminates after a maximum number of iterations (i.e., when $k>K$) or when parameters have converged (i.e., $\|\theta_{k+1}-\theta_{k}\|_{2}\leq\epsilon$). In our experiments, $K=500$ and $\epsilon=10^{-4}$. At each step, also note that we project the discount factor $\gamma$ onto the set $[0,\infty)^{N}$ to ensure feasibility.
1 Hyper-parameters: Learning rate $\alpha$
@@ 5 $k\leftarrow 0$
7       $v_{k}\leftarrow\operatorname{solveInnerMCP}(\theta_{k},\gamma_{k})$triangleright Equation 10
8       $\nabla_{(\theta_{k},\gamma_{k})}\mathcal{P}\leftarrow\operatorname{calcGradient}(v_{k},\theta_{k},\gamma_{k})$triangleright Equation 11Equation 12
@@ VI-A Example Game - Crosswalk
Our first experiment involves $N=2$ pedestrians navigating a crosswalk at an intersection on the UT Austin campus at W Dean Keaton St. and Whitis Ave. This intersection becomes safe for all pedestrian crossing at once, encouraging students to cross diagonally and interact with one another. We simulate two short-sighted students who begin at the top right and left corners, respectively, and their goals are to reach the bottom left and right corners, respectively.
We model the students as wishing to minimize the distance to their goal and their control effort (they are tired), such that the $i^{\mathrm{th}}$ agent’s cost can be written as:
$J^{i}_{\mathrm{stud}}=\sum_{t=0}^{T}(\gamma^{i})^{t}\bigr{[}w_{\textrm{goal}}^{i}\|p^{i}_{t}-\theta^{i}_{\textrm{goal}}\|^{2}_{2}+w_{\textrm{ctrl}}^{i}\|u^{i}_{t}\|^{2}_{2}].$ (13)
[…]
$x_{t+1}^{i}=\begin{bmatrix}1&0&\Delta&0\\ 0&1&0&\Delta\\ 0&0&1&0\\ 0&0&0&1\end{bmatrix}\underbrace{\begin{bmatrix}p_{x,t}^{i}\\ p_{y,t}^{i}\\ v_{x,t}^{i}\\ v_{y,t}^{i}\end{bmatrix}}_{x_{t}^{i}}+\begin{bmatrix}\nicefrac{{\Delta^{2}}}{{2}}&0\\ 0&\nicefrac{{\Delta^{2}}}{{2}}\\ \Delta&0\\ 0&\Delta\end{bmatrix}\underbrace{\begin{bmatrix}a_{x,t}^{i}\\ a_{y,t}^{i}\end{bmatrix}}_{u_{t}^{i}}$
@@
where $d_{\textrm{min}}$ denotes the minimum allowed distance between the agents and $\delta>1$ is a safety margin. In our testing, we found that Algorithm 1 converged more reliably when these constraints are implemented as large penalties in each agent’s cost function. Therefore, we model $i^{\mathrm{th}}$ agent’s objective function in the inverse game as $J^{i}_{\mathrm{ped}}=J^{i}_{\mathrm{stud}}+J^{i}_{\mathrm{coll}}$, where
$J^{i}_{\mathrm{coll}}=\sum_{t=0}^{T}(\gamma^{i})^{t}w_{\textrm{coll}}^{i}\max(0,\delta\cdot d_{\textrm{min}}-\|p^{i}_{t}-p^{\neg i}_{t}\|^{2}_{2}\bigl{)},$ (15)
@@ VI-B Example Game - Intersection
Figure 2: Intersection from the InD dataset [inDdataset], overlaid with trajectories for all four agents generated in a receding horizon fashion using our method and the baseline. Note that we reduced the opacity of the complete foresighted rollout after $t$ = 15 for visual clarity.
Our second experiment involves $N=4$ agents navigating an intersection: two pedestrians crossing a crosswalk, a car attempting to turn while the pedestrians are crossing, and a car attempting to go straight while behind the turning car. We employ real-world data of this interaction from the Intersection Drone (InD) dataset, captured by leveLXData via a drone hovering over the intersection found in Aachen, Germany, seen in Figure 2 [inDdataset]. Specifically, the interaction takes place over 162 frames at 25 frames per second. The data was downsampled by a factor of 6, resulting in a game horizon of $T=28$.
@@
We model the cars’ physical dynamics as bicycles with bounds on the control inputs. Explicitly, the game’s state at time $t$ is composed as $x_{t}=[x^{1\top}_{t},x^{2\top}_{t}]^{\top}$, and for the $i^{\mathrm{th}}$ agent:
${x}_{t+1}^{i}=\underbrace{\begin{bmatrix}p_{x,t}^{i}\\ p_{y,t}^{i}\\ v_{t}^{i}\\ \psi_{t}^{i}\end{bmatrix}}_{x_{t}^{i}}+\Delta\begin{bmatrix}v^{i}_{t}\cos(\psi_{t}^{i})\\ v^{i}_{t}\sin(\psi_{t}^{i})\\ a_{t}^{i}\\ v_{t}^{i}\tan(\phi_{t}^{i})/l\end{bmatrix}$
@@
Depending on whether the $i^{\mathrm{th}}$ agent is a car or pedestrian, we model the agent’s objective function slightly differently. For the pedestrians, their cost is exactly $J^{i}_{\mathrm{ped}}$. Meanwhile, the cars aim to be in the center of their lanes and, thus, have an additional term that minimizes the distance to their lane centers. We model the $i^{\mathrm{th}}$ agent’s lane center as a set, $\mathcal{C}^{i}=\{(c_{x,t}^{i},c_{y,t}^{i})\}_{t=0}^{T}$, generated from a polynomial curve fit to lane center points. We then can write the cost as:
$J^{i}_{\mathrm{car}}=J^{i}_{\mathrm{ped}}+\sum_{t=0}^{T}(\gamma^{i})^{t}w_{\textrm{lane}}^{i}\|p^{i}_{t}-c^{i}_{t}\|^{2}_{2},$ (16)
@@
VI-C Experimental Setup
Figure 3: Heatmap of $\mathcal{P}$ from 1 in a fully observable (on the left) and partially observable setting (on the right). In both cases, both player’s $\gamma$ are varied while holding $\theta$ constant at the ground truth values and the resulting cost is scaled to be in between $[0,1]$.
@@ VI-C Experimental Setup
We aim to show that our method predicts trajectories more accurately than a baseline method, which solves the same inverse problem but presumes agents have a discount factor of $1$. [^1] For more details on our implementation, refer to our GitHub repository. Thus, the baseline estimator has fewer hidden parameters to estimate than our method.
To ensure a fair comparison, we use the same MiCP back-end [dirkse1995path] to satisfy the first-order necessary conditions of the GOLNE at line 1 of Algorithm 1. Additionally, we set the default learning rate to $\alpha=10^{-3}$ for both methods. As input, both methods receive the same observation sequence at every simulation, as well as the same initial $\theta$; our method receives a uniformly random $\gamma\in[0,1]$ as an initial estimate.
[…]
VI-D Monte Carlo Results - Crosswalk
(a) With only position measurements
(b) With full state measurements
Figure 4: Algorithm 1 reliably improves upon the baseline, indicating that modeling and estimating agents’ foresightedness leads to better trajectory predictions. These results hold in settings with both partial state observations and full state observations. Solid/dotted lines denote means and the opaque band indicates standard deviation.
@@ VI-D Monte Carlo Results - Crosswalk
Figure 1 shows a representative trial for our running example at a simulated crosswalk. It shows the two players’ initial positions, their respective ground-truth goals, their inferred goals for the baseline and our method, the observed trajectory being recovered, and their current and planned trajectories at a snapshot in time. Note that our method results in a more visually similar trajectory in comparison to the observations. In addition, according to the bar graph, our method accurately recovers the discount factor within $0.3\%$.
@@
Figure 3 shows a heatmap visualizing the objective function for 1 in which $\theta$ is fixed to the ground truth values, in both fully- and partially-observed settings. Here, the ground truth values of the agents’ discount factors are $\gamma^{1}=\gamma^{2}=0.6$. As shown, in both settings, the inverse game’s objective function has a global minimum at this point. However, we also observe that there is a narrow, diagonally-shaped low-cost region: this implies that the equilibrium of the game is relatively insensitive to agents’ discount factors as long as both agents have similar discount factors (and both are greater than $\approx 0.5$). While it is intuitive, perhaps, that the relative performance of the agents should exhibit this kind of symmetry, it is not immediately clear that equilibria of the game should do so.
Nevertheless, this observation motivates us to introduce the following regularization scheme for the inverse game in 1.
Problem 2.
Regularized Inverse Game Problem: Given a sequence of observations y, find parameters $\hat{\theta}$ which solve
$\min_{\textbf{x},\textbf{u},\theta,\gamma}~{}$ $\mathcal{P}\left(\textbf{x}(\theta,\gamma)\right)+c_{\gamma}\|1-\gamma\|_{2}^{2}$ (17a)
[…]
In 2, we have introduced the term $\|1-\gamma\|^{2}_{2}$, which introduces a slight bias toward higher discount factors. Note, however, that this essentially encodes a prior belief that agents are more farsighted; it is straightforward to instead penalize $\|\gamma\|_{2}^{2}$ and encode the belief that agents are more shortsighted. In the following Monte Carlo results, we report results for $c_{\gamma}=10^{-3}$.
@@
Figure 4 shows the mean trajectory estimation error as a function of the observation noise standard deviation for Algorithm 1 and the $\gamma=1$ baseline. Our method substantially outperforms the baseline, exhibiting about a $30\%$ more accurate recovered trajectory in the partially observable case, and a decreasing improvement with more noise in the fully observable case. Interestingly, we can see that the mean error is only marginally sensitive to noise for both methods. Rather, the most noticeable effect of noise is an increase in the standard deviation of the errors. From these results, we conclude that our foresighted formulation can better predict the behavior of short-sighted agents in comparison to a state-of-the-art inverse game baseline [liu2023learningplaytrajectorygames].
VI-E Monte Carlo Results - Intersection
(a) With only position measurements
(b) With full state measurements
Figure 5: Similarly to Figure 4, we show Algorithm 1 continues to improve performance on the baseline, even with real-world data. These results hold in settings with both (a) partial state observations and (b) full state observations. Solid lines show error means, while the opaque bands show the standard deviations. Individual experiment results are plotted as dots.

### v3

@@ V-B Optimizing Parameters with Gradient Descent
The only non-trivial term in Equation 7 is $\nabla_{(\theta,\gamma)}v$. Next, we show how to take directional derivatives of MiCPs at a solution $v^{*}$ where strictly complementarity holds.
First, we must consider the complementarity constraints on $z$ and $h(\cdot)$. To this end, we construct an index set $\mathcal{I}$ which records all inactive inequality constraint dimensions in $h(v;\theta,\gamma)$. Indexing $F$ at elements of this set yields a vector $F]_{\mathcal{I}}$. From Section III-C, we know that these inactive inequalities $F]_{\mathcal{I}}$ are strictly positive, and the Lagrange multipliers associated with these constraints are exactly 0. Then, presuming the continuity of $F$, small changes in $(\theta,\gamma)$ preserve the positivity of elements of $F]_{\mathcal{I}}$ and force the corresponding Lagrange multipliers in $v]_{\mathcal{I}}$ to remain 0. Thus, we find that $\nabla_{(\theta,\gamma)}[v]_{\mathcal{I}}=0$.
Consider the remaining constraints from Equation 6, which must be active due to strict complementarity; denote the indices corresponding to these constraint as $\mathcal{S}$. We can use the implicit function theorem and stationarity of $F$ with respect to $v$ to write:
$0$ $=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}=\nabla_{(\theta,\gamma)}[F]_{\mathcal{S}}+(\nabla_{v}[F]_{\mathcal{S}})(\nabla_{(\theta,\gamma)}v)$ (8a)
@@
Then, when $\nabla_{v}[F]_{\mathcal{S}}$ is invertible, we can find exact values of $\nabla_{(\theta,\gamma)}v$. For a more complete treatment, including a discussion of weak complementarity, readers are encouraged to consult [47, 42, 18].
Thus equipped, we iteratively update our estimate of $(\theta,\gamma)$ according to the gradient of Equation 4a—incorporating the aforementioned implicit derivatives as needed. Algorithm 1 terminates after a maximum number of iterations (i.e., when $k>K$) or when parameters have converged (i.e., $\|\theta_{k+1}-\theta_{k}\|_{2}\leq\epsilon$). In our experiments, $K=500$ and $\epsilon=10^{-4}$. At each step, also note that we project the discount factor $\gamma$ onto the set $0,\infty)^{N}$ to ensure feasibility.
1Hyper-parameters: Learning rate $\alpha$
@@ 5 $k\leftarrow 0$
7 $v_{k}\leftarrow\operatorname{solveInnerMCP}(\theta_{k},\gamma_{k})$ $\triangleright$ Section V-A
8 $\nabla_{(\theta_{k},\gamma_{k})}\mathcal{P}\leftarrow{{\color[rgb]{0,0,0}{\operatorname{calcGrad}}}}(v_{k},\theta_{k},\gamma_{k})$ $\triangleright$ Section V-B
@@ VI-A Example Game - Crosswalk
Our first experiment involves $N=2$ pedestrians navigating a crosswalk at an intersection on the UT Austin campus at W Dean Keaton St. and Whitis Ave. This intersection becomes safe for all pedestrian crossing at once, encouraging students to cross diagonally and interact with one another. We simulate two short-sighted delivery robots who begin at the top right and left corners, respectively, and their goals are to reach the bottom left and right corners, respectively.
We model agents as delivery robots wishing to minimize the distance to their goal and their control effort, such that the $i^{\mathrm{th}}$ agent’s cost can be written as:
$J^{i}_{\mathrm{stud}}=\sum_{t=0}^{T}(\gamma^{i})^{t}~\bigr[w_{\textrm{goal}}^{i}~\|p^{i}_{t}-\theta^{i}_{\textrm{goal}}\|^{2}_{2}+w_{\textrm{ctrl}}^{i}~\|u^{i}_{t}\|^{2}_{2}].$ (9)
For brevity, we define $p^{i}_{t}\in\mathbb{R}^{2}$ to denote agent $i$’s position at time $t$, weigh the goal and control cost terms with $w^{i}_{\textrm{goal}}$ and $w^{i}_{\textrm{ctrl}}$, and parameterize the game with the $i^{\mathrm{th}}$ agent’s target position $\theta^{i}_{\textrm{goal}}\in\mathbb{R}^{2}$. We model the robots as point masses with bounds on the velocity and control inputs, assume a time discretization of $\Delta=$0.1\text{\,}\mathrm{s}$\$, and play the game over $T=25$ time steps. Explicitly, the game’s state at time $t$ is composed as $x_{t}=[x^{1\top}_{t},x^{2\top}_{t}]^{\top}$, and each agent’s physical dynamics are modeled with a standard double integrator linear system.
@@
where $d_{\textrm{min}}$ denotes the minimum allowed distance between the agents and $\delta>1$ is a safety margin. However, by moving this constraint to each agent’s objective, we can avoid assigning a new Lagrange multiplier to each pair of agents. This avoids quadratic scaling in the number of MiCP variables with respect to number of agents. Therefore, we model $i^{\mathrm{th}}$ agent’s objective function in the inverse game as $J^{i}_{\mathrm{ped}}=J^{i}_{\mathrm{stud}}+J^{i}_{\mathrm{coll}}$, where
$J^{i}_{\mathrm{coll}}=\sum_{t=0}^{T}(\gamma^{i})^{t}~w_{\textrm{coll}}^{i}~\max(0,\delta\cdot d_{\textrm{min}}-\|p^{i}_{t}-p^{\neg i}_{t}\|^{2}_{2}\bigl),$ (11)
@@ VI-B Example Game - Intersection
Figure 2: Intersection from the InD dataset [48], overlaid with trajectories for all four agents generated in a receding horizon fashion using our method and the baseline. Note that we reduced the opacity of the complete foresighted rollout after $t$ = 15 for visual clarity.
Our second experiment involves $N=4$ agents navigating an intersection: two pedestrians crossing a crosswalk, a car attempting to turn while the pedestrians are crossing, and a car attempting to go straight while behind the turning car. We employ real-world data of this interaction from the Intersection Drone (InD) dataset, captured by leveLXData via a drone hovering over the intersection found in Aachen, Germany, seen in Figure 2 [48]. Specifically, the interaction takes place over 162 frames at 25 frames per second. The data was downsampled by a factor of 6, resulting in a game horizon of $T=28$.
@@
We model the cars’ physical dynamics as bicycles with bounds on the control inputs. Explicitly, the game’s state at time $t$ is composed as $x_{t}=[x^{1\top}_{t},x^{2\top}_{t}]^{\top}$ in which $x_{t}^{i}=[p_{x,t}^{i},p_{y,t}^{i},v_{t}^{i},\psi_{t}^{i}]^{\top}$, and for the $i^{\mathrm{th}}$ agent:
${x}_{t+1}^{i}=$ $[p_{x,t}^{i},~p_{y,t}^{i},~v_{t}^{i},~\psi_{t}^{i}]^{\top}+$
$\Delta$ $[v^{i}_{t}\cos(\psi_{t}^{i}),~v^{i}_{t}\sin(\psi_{t}^{i}),~a_{t}^{i},~v_{t}^{i}\tan(\phi_{t}^{i})/l]^{\top}$
@@
Depending on whether the $i^{\mathrm{th}}$ agent is a car or pedestrian, we model the agent’s objective function slightly differently. For the pedestrians, their cost is exactly $J^{i}_{\mathrm{ped}}$ while cars also minimize the distance to their lane centers. We model the $i^{\mathrm{th}}$ agent’s lane center as a set, $\mathcal{C}^{i}=\{(c_{x,t}^{i},c_{y,t}^{i})\}_{t=0}^{T}$ generated from a polynomial curve fit to lane center points. We then can write the cost as:
$J^{i}_{\mathrm{car}}=J^{i}_{\mathrm{ped}}+\sum_{t=0}^{T}(\gamma^{i})^{t}~w_{\textrm{lane}}^{i}~\|p^{i}_{t}-c^{i}_{t}\|^{2}_{2},$ (12)
@@
VI-C Example Game - Waymax Simulation
Figure 3: Snapshots from receding horizon inference and planning with the Waymax simulator, for a game involving the blue, green, yellow, and red agents. The ego robot (blue) employs our method, while the other cars are controlled by the Waymax simulator.
Our last experiment involves $N=4$ cars navigating an intersection, one of which is a robot attempting to make a left turn (cf. Figure 3). We employ the Waymax simulator, which has been built atop the Waymo Open Motion Dataset [49], to simulate the humans’ responses to the ego’s actions. The ego robot infers the humans’ objectives and foresightedness and plans to reach its goal safely in a receding horizon setup. Specifically, the scenario takes place over a horizon of $T=50$ steps with a time discretization of $\Delta=$0.1\text{\,}\mathrm{s}$\$ and a planning horizon $T_{\mathrm{plan}}=11$ steps.
Like in Section VI-B, we model the cars’ physical dynamics as bicycles with bounds on the control inputs. However, we modify the costs from the previous examples by: 1) simplifying the lane center term to only consider $p_{x,t}^{i}$ as the road is a relatively straight line after the turn, and 2) adding a quadratic target velocity tracking cost. Lastly, it should be noted that the game is constructed such that the ego agent has no parameters in its own cost or constraints, so that it is only inferring the parameters of the simulated agents.
VI-D Experimental Setup
@@
For our first two examples, we aim to show that our method predicts trajectories more accurately than a baseline method, which solves the same inverse problem but presumes agents have a discount factor of 11.111Full implementation details can be found at https://github.com/cadearmstrxng/InverseGameDiscountFactor.jl. Thus, the baseline estimator has fewer hidden parameters to estimate than our method.
To ensure a fair comparison, we use the same MiCP back-end [43] to satisfy the first-order necessary conditions of the GOLNE at line 1 of Algorithm 1, the same observation sequence, and the same initial guess—our method receives a uniformly random $\gamma\in[0,1]$ as an initial estimate. To compare the two methods, we calculate their respective error as $\textrm{error}=\|\textbf{x}-\textbf{x}_{\textrm{ground-truth}}\|^{2}_{2}$, enabling us to determine how well the methods predicted the agents’ trajectories. We gather these results with a Monte Carlo study done over a range of observation noise covariances as previously discussed. For full state experiments, the mean, $h(x_{t})$, is the identity function. For partial state experiments, $h(x_{t})=[p_{t}^{i\top}~\forall i\in N]^{\top}$. The covariance is given by identity matrices scaled by 50 evenly spaced constants between 0 and $0.01\text{\,}{\mathrm{m}}^{2}$. For each covariance matrix, we generated 50 noisy observation sequences given to our method and the baseline method, resulting in $2500$ observation sequences and parameter estimates. Results are shown in Figure 5 and Figure 6, and discussed below.
Figure 4: Heatmap of $\mathcal{P}$ from 1 for the crosswalk experiment for the fully observable (left) and partially observable (right) cases. Both players’ $\gamma$ are varied while holding $\theta$ constant at the ground truth values and the resulting cost is scaled to be in between $[0,1]$.
For the Waymax example, we aim to show our method can be used in a robotics context safely and efficiently by employing it in a receding horizon planner. At each time $t$, the ego robot infers other agents’ cost parameters from the past 10 observations and plans over the next 10 time steps. The action used at $t$ is the first action taken in the planning trajectory. To evaluate safety and efficiency, we simulate 50 random sequences of observation noise with standard deviation $0.1\text{\,}\mathrm{m}$ and plot the ego robot’s distance to each other robot and to its goal position over the time horizon, shown in Figure 7. Results are qualitatively similar at lower noise levels.
VI-E Monte Carlo Results - Crosswalk
@@
Figure 1 shows a representative trial for our running example at a simulated crosswalk. It shows the two players’ observations, initial positions, ground-truth goals, inferred goals for the baseline and our method, and current and planned trajectories at a snapshot in time. Note that our method results in a trajectory closer to the observations in L2 distance than the baseline. In addition, according to the bar graph, our method accurately recovers the discount factor within $0.3\%$.
@@
Figure 4 shows a heatmap visualizing the objective function for 1 in which $\theta$ is fixed to the ground truth values, in both fully- and partially-observed settings. The diagonally-shaped low-cost region implies that the equilibrium of the game is relatively insensitive to the agents’ absolute discount factors as long as both agents have similar discount factors (and both are greater than $\approx 0.5$). We expect this low-cost region to exist because of the game’s highly symmetric structure, motivating us to introduce a regularization scheme for the inverse game in 1 to encourage better solver performance. To do so, we added the term $c_{\gamma}\|1-\gamma\|_{2}^{2}$ to Equation 4a, which encodes a prior belief that agents are more farsighted; it is straightforward to instead penalize $\|\gamma\|_{2}^{2}$ and encode the belief that agents are more shortsighted. In the following Monte Carlo results, we report results for $c_{\gamma}=10^{-3}$. For clarity, we do not regularize in the other, asymmetric experiments.
@@
Figure 5 shows the mean trajectory estimation error as a function of the observation noise standard deviation for Algorithm 1 and the $\gamma=1$ baseline. Our method substantially outperforms the baseline in both cases. Interestingly, we can see that the mean error is only marginally sensitive to noise for both methods. Rather, the most noticeable effect of noise is an increase in the standard deviation of the errors. From these results, we conclude that our foresighted formulation can better predict the behavior of short-sighted agents in comparison to a state-of-the-art inverse game baseline [18].
(a) Only position measurements
(b) Full state measurements
Figure 5: Algorithm 1 reliably improves predicted trajectory error in the crosswalk experiment, for partial and full state observations. Solid/dotted lines denote means and the opaque band indicates standard deviation.
VI-F Monte Carlo Results - InD Intersection